OpenStudy (ethan01):

Given: Q = 7m + 3n, R = 11 - 2m, S = n + 5, and T = -m - 3n + 8. Simplify Q + S - T.

OpenStudy (jtug6):

Ok so you know that Q is 7m + 3n, right? And r is 11 - 2m just as S = n+ 5 and T = -m -3n + 8. And you want Q + S - T so plug in what Q, S, and -T equal from what we're given. (7m + 3n) + (n + 5) - (-m -3n + 8)

OpenStudy (jtug6):

So now just combine like terms. Also note that youre subtracting T, or adding negative T. So, carry the negative through T and then add that ^_^

OpenStudy (ethan01):

ok thanks

OpenStudy (jtug6):

np buddy